OpenStudy (anonymous):

y/-6+5=9 . solve for y

OpenStudy (anonymous):

First you would take the -6 and the 5 which would equal to -1, this gives you y/-1 then you multiple the -1 over the the 9, which gives you y=-9. I believe.

OpenStudy (ness9630):

First what you do is move that 5 over, giving you -y/6=4 Now just multiply everything by -6 and that will get y by itself

OpenStudy (ness9630):

Is this your equation? \[\frac{y}{-6}+5=9\]

OpenStudy (anonymous):

yes @Ness9630

OpenStudy (anonymous):

I thought the whole thing was under the y.

OpenStudy (ness9630):

Then follow my steps

OpenStudy (ness9630):

So you are left with: \[-\frac{y}{6}=5\] Just multiply by -6

OpenStudy (anonymous):

wouldnt it be -24?

OpenStudy (ness9630):

Oh, excuse me that should be a 4, not a 5. And yes, it would be -24 Good job!